# Simultaneous symmetry breaking in spontaneous Floquet states: Floquet-Nambu-Goldstone modes, Floquet thermodynamics, and the time operator

### v2

@@ VI.1 $(t,\phi)$ formalism
In order to correctly derive the dynamics of the temporal FNG mode, we develop the $(t,\phi)$ formalism, denoted in this way due to its analogy with the more usual $(t,t^{\prime})$ formalism Peskin and Moiseyev (1993); Heimsoth et al. (2012), where $\phi\in[0,2\pi)$ is the angular variable associated to the periodic motion of the spontaneous Floquet state, $\phi=\omega t^{\prime}$, and $t$ is a proper time that accounts for the nonperiodic dynamics. As we will see, the use of $\phi$ instead of $t^{\prime}$ is critical. In the $(t,\phi)$ formalism, we add an extra dimension to the field, $\Psi=\Psi(\mathbf{x},\phi,t)$, and postulate that the dynamics is described by the Lagrangian
$L=\frac{1}{2\pi}\int^{2\pi}_{0}\mathrm{d}\phi\left(\int\mathrm{d}\mathbf{x}~{}[i\Psi^{*}\partial_{t}\Psi+i\Psi^{*}\omega\partial_{\phi}\Psi]-K\right),$ (127)
where $K$ does not explicitly depend on $\phi$ as the Hamiltonian is time-independent. The case of a conventional, driven Floquet system is simply accounted by a $\phi$-dependent periodic Hamiltonian, $H(\phi+2\pi)=H(\phi)$, with $H(\phi)\equiv H(\omega t)$.
The action is the time integral of the Lagrangian (127), to which we apply Noether’s Theorem in order to obtain the conserved charges associated to the continuous symmetries. The extra variable $\phi$ simply adds an angular average to Eq. (12). In particular, the usual charges $\mathbf{Q}$ are readily retrieved as
@@
the scalar product in the extended Hilbert space. Since they are independent of $\phi$ in virtue of the original Lagrangian (21), $\mathbf{Q}(\phi)=\mathbf{Q}$, we recover the same values for the charges as before.
@@
with $T_{\phi}=i\partial_{\phi}$ the generator of phase translations. The operator $T_{\phi}$ is Hermitian under the extended scalar product (VI.1), and commutes with the rest of generators, $[T_{\phi},\mathbf{T}]=0$.
The invariance under time translations $t\to t+t_{0}$ of the Lagrangian (127) results in the conservation of the generalized Floquet-Gibbs energy $\Lambda$,
$\Lambda=K-\omega F=H-\boldsymbol{\lambda}\cdot\mathbf{Q}-\omega F=H-\lambda^{\alpha}Q_{\alpha},$ (131)
which is the extension of the generalized Gibbs energy $K$ for Floquet states. Its expression allows us to identify $\lambda^{\phi}=\omega$ as the Lagrange multiplier of the Floquet charge. We note that both $F,\Lambda$ need to include an angular average to be truly time-independent, in contrast to the conventional charges $\mathbf{Q}$. However, while $\Lambda$ is conserved for both driven and spontaneous Floquet states, the conservation of the Floquet charge $F$ is a genuine feature of a spontaneous Floquet state since the Hamiltonian is $\phi$-dependent for driven Floquet systems.
@@ VI.2 Spontaneous Floquet states in the generalized Gibbs ensemble: Floquet thermodynamics
The Floquet-Gibbs energy $\Lambda$ is now the thermodynamic potential which is extremized by stationary solutions $\Psi_{0}(\mathbf{x},\phi)$, satisfying the Floquet-GPG (FGPG) equation
@@
When we insert the FGPG wavefunction in the actual time-dependent GP equation (20) by taking $\Psi(\mathbf{x},t=0)=\Psi_{0}(\mathbf{x},\phi)$, we find
$\Psi(\mathbf{x},t)=e^{-i\lambda^{\alpha}T_{\alpha}t}\Psi_{0}(\mathbf{x},\phi)=e^{-i\boldsymbol{\lambda}\cdot\mathbf{T}t}\Psi_{0}(\mathbf{x},\phi+\omega t).$ (133)
Thus, Floquet states are described by stationary FGPG solutions with the phase $\phi$ fixed to some value $\phi_{0}$, which behaves as a global time origin $t_{0}=\phi_{0}/\omega$. When the system is driven, this time origin is locked by the actual time-dependence of the Hamiltonian, $H(t)=H(\phi_{0}+\omega t)$, which we can conventionally set to $\phi_{0}=0$. However, for spontaneous Floquet states, $\phi_{0}$ is a free parameter, reflecting the time-independence of the original Hamiltonian. On the other hand, as expected from thermodynamic considerations, $\omega$ plays the predicted role for the Lagrange multiplier associated to the Floquet charge, acting as the velocity of phase translations. In fact, the definition modulo $\omega$ of the chemical potential $\mu$ for Floquet states can be also explained from Thermodynamics as follows. The transformation $\Psi_{0}(x,\phi)\to\Psi_{0}(x,\phi)e^{-in\phi}$, $n\in\mathbb{Z}$, also yields a solution of the FGPG equation. This leads to $F\to F+nN$, so $\Lambda\to\Lambda-n\omega N$, and hence $\lambda^{\theta}=\mu$ changes as $\mu\to\mu+n\omega$.
@@
we obtain the generalization of the first principle of Thermodynamics
@@
and the Gibbs-Duhem relation for Floquet states, Eq. (32).
The upshot of the above discussion is that Floquet states allow for a thermodynamical description completely analogous to that of stationary states, which we will label as Floquet thermodynamics. Indeed, the first principle of Thermodynamics for Floquet states (135) suggests a fundamental distinction between spontaneous and driven Floquet systems: spontaneous systems conserve the charge $F$ and the frequency $\omega$ is determined by the equation of state, while driven Floquet systems externally impose the frequency $\omega$ so then $F$ is not conserved. Thus, we can regard spontaneous Floquet states as “isofloquetic”, where the macroscopic Floquet charge is conserved, in analogy with isentropic or isochoric systems, where the entropy or the volume is conserved. Conversely, conventional Floquet systems are “isoperiodic”, where the external driving fixes the oscillation period, in analogy with isothermal or isobaric systems, where the external environment fixes the temperature or the pressure. As a result, the natural thermodynamic potential for isoperiodic systems is the “Floquet enthalpy”
$I\equiv E-\omega F=\Lambda-\boldsymbol{\lambda}\cdot\mathbf{Q},$ (136)
which is indeed conserved for driven systems, and satisfies
@@
VI.3 Quantization of the Floquet-Nambu-Goldstone modes
The $(t,\phi)$ formalism emerges in the quantum theory when considering once more the time-independent Heisenber equation (62). We now perform a unitary transformation to the field operator along the lines of Eq. (133),
$\hat{\Psi}(\mathbf{x},t)=\left.e^{-i\lambda^{\alpha}T_{\alpha}t}\hat{\Psi}(\mathbf{x},\phi,t)\right|_{\phi=\phi_{0}},$ (138)
obtaining a $\phi$-dependent field operator. After expanding around the spontaneous Floquet state, $\hat{\Psi}(\mathbf{x},\phi,t)=\Psi_{0}(\mathbf{x},\phi)+\hat{\varphi}(\mathbf{x},\phi,t)$, we obtain at linear order
@@ VI.3 Quantization of the Floquet-Nambu-Goldstone modes
Once the dynamics in the $(t,\phi)$ formalism is derived, by taking $\phi=\phi_{0}$ we recover the GPG wavefunction $\Psi_{0}(x,t)=\Psi_{0}(x,\phi_{0}+\omega t)$ and its quantum fluctuations $\hat{\varphi}(\mathbf{x},t)=\hat{\varphi}(\mathbf{x},\phi_{0}+\omega t,t)$. Notice that $\phi_{0}$ can be freely chosen for a spontaneous Floquet state but, after locked by the symmetry breaking, it is also fixed for $\hat{\varphi}$.
@@
The Goldstone-Gibbs modes $z_{a}(\mathbf{x},\phi)$ are given here in terms of $\partial_{a}\Psi_{0}(\mathbf{x},\phi)$, while the Bogoliubov sector $z_{\varepsilon,\nu}(\mathbf{x},\phi)$ is obtained by the adaptation of the Floquet modes of Eq. (121). In order to have a complete orthonormal basis in the extended Hilbert space, we promote each mode $z_{n}(\mathbf{x},\phi)$ to a whole set
$z_{n,m}(\mathbf{x},\phi)\equiv z_{n}(\mathbf{x},\phi)e^{im\phi},~{}m\in\mathbb{Z},$ (141)
@@
$[M_{0}(\phi)-i\omega\partial_{\phi}]z_{\alpha,m}=m\omega z_{\alpha,m}$
$[M_{0}(\phi)-i\omega\partial_{\phi}]z_{A,m}=m\omega z_{A,m}+i\partial_{A}\lambda^{\beta}z_{\beta,m}$ (142)
$[M_{0}(\phi)-i\omega\partial_{\phi}]z_{\nu,\varepsilon,m}=[\varepsilon+m\omega]z_{\nu,\varepsilon,m},$
@@
$(z_{\nu,\varepsilon,m}||z_{\nu^{\prime},\varepsilon^{\prime},m^{\prime}})$ $=$ $\delta_{\nu\nu^{\prime}}\delta_{\varepsilon\varepsilon^{\prime}}\delta_{mm^{\prime}},$
$(z_{a,m}||z_{\nu,\varepsilon,m^{\prime}}))$ $=$ $0,$ (143)
@@
Their quantum amplitude is now obtained by $\hat{\gamma}_{n,m}(t)=(z_{n,m}||\hat{\Phi}(t))$. By using the Fourier relations
@@
$\hat{\Phi}(\mathbf{x},\phi,t)$ $=$ $\sum_{m=-\infty}^{\infty}\hat{X}_{m}^{\alpha}(t)z_{\alpha,m}(\mathbf{x},\phi)+\hat{P}_{m}^{A}(t)z_{A,m}(\mathbf{x},\phi)+\sum_{m=-\infty}^{\infty}\sum_{\nu,\varepsilon}\hat{\gamma}_{\nu,\varepsilon,m}(t)z_{\nu,\varepsilon,m}(\mathbf{x},\phi)+\hat{\gamma}^{\dagger}_{\nu,\varepsilon.m}(t)\bar{z}_{\nu,\varepsilon,m}(\mathbf{x},\phi)$ (145)
$=$ $\hat{X}^{\alpha}(\phi,t)z_{\alpha}(\mathbf{x},\phi)+\hat{P}^{A}(\phi,t)z_{A}(\mathbf{x},\phi)+\sum_{\nu,\varepsilon}\hat{\gamma}_{\nu,\varepsilon}(\phi,t)z_{\nu,\varepsilon}(\mathbf{x},\phi)+\hat{\gamma}^{\dagger}_{\nu,\varepsilon}(\phi,t)\bar{z}_{\nu,\varepsilon}(\mathbf{x},\phi).$
The dynamics for $\hat{\Phi}(\mathbf{x},\phi,t)$ is readily obtained from
$\hat{\gamma}_{\nu,\varepsilon,m}(t)=\hat{\gamma}_{\nu,\varepsilon,m}e^{-i[\varepsilon+m\omega]t},~{}~{}~{}\hat{P}_{m}^{A}(t)=\hat{P}_{m}^{A}e^{-im\omega t},~{}~{}~{}\hat{X}_{m}^{\alpha}(t)=[\hat{X}_{m}^{\alpha}+\partial_{A}\lambda_{B}\hat{P}_{m}^{B}t]e^{-im\omega t}.$ (146)
When setting $\phi=\phi_{0}+\omega t$, after noticing that $z_{n}(\mathbf{x},t)=z_{n}(\mathbf{x},\phi_{0}+\omega t)$ and $\hat{\gamma}_{n}(t)=\hat{\gamma}_{n}(\phi_{0}+\omega t,t)$, we recover the expected time evolution for the field fluctuations:
$\hat{\Phi}(\mathbf{x},t)=\hat{\Phi}(\mathbf{x},\phi_{0}+\omega t,t)=[\hat{X}^{\alpha}+\partial_{A}\lambda_{B}\hat{P}^{B}t]z_{\alpha}(\mathbf{x},t)+\hat{P}^{A}z_{A}(\mathbf{x},t)+\sum_{\nu,\varepsilon}\hat{\gamma}_{\nu,\varepsilon}z_{\nu,\varepsilon}(\mathbf{x},t)e^{-i\varepsilon t}+\hat{\gamma}^{\dagger}_{\nu,\varepsilon}\bar{z}_{\nu,\varepsilon}(\mathbf{x},t)e^{i\varepsilon t}.$ (147)
We note that the above expansion is also valid for driven Floquet systems by excluding the Goldstone-Gibbs pair $\{z_{\phi},z_{F}\}$ and using $\lambda_{A}=\partial_{A}I$ instead of $\lambda_{A}=\partial_{A}E$. On the other hand, the crucial difference with respect to Eq. (VI) is that the Gibbs modes there encode a non-trivial trivial dependence in $z_{\phi}$ since
@@
When this dependence is taken into account, as well as the thermodynamic relation between $K$ and $F$, after some tedious algebra one arrives at the correct expansion (147) from the original Floquet modes of Eqs. (120), (121).

### v3

@@ VI.1 $(t,\phi)$ formalism
In order to correctly derive the dynamics of the temporal FNG mode, we develop the $(t,\phi)$ formalism, denoted in this way due to its analogy with the more usual $(t,t^{\prime})$ formalism Peskin and Moiseyev (1993); Heimsoth et al. (2012). Here, $\phi\in[0,2\pi)$ is the angular variable associated to the periodic motion of the spontaneous Floquet state, $\phi=\omega t^{\prime}$, and $t$ is a proper time that accounts for the nonperiodic dynamics. As we will see, the use of $\phi$ instead of $t^{\prime}$ is critical.
In the $(t,\phi)$ formalism, we add an extra dimension to the field, $\Psi=\Psi(\mathbf{x},\phi,t)$, whose dynamics is described by the Lagrangian
$L=\frac{1}{2\pi}\int^{2\pi}_{0}\mathrm{d}\phi\left(\int\mathrm{d}\mathbf{x}~{}[i\Psi^{*}\partial_{t}\Psi+i\Psi^{*}\omega\partial_{\phi}\Psi]-H\right),$ (127)
where $H$ does not explicitly depend on $\phi$ as the Hamiltonian is time-independent. The case of a conventional, driven Floquet system is simply accounted by a $\phi$-dependent periodic Hamiltonian, $H(\phi+2\pi)=H(\phi)$.
The resulting equation of motion for $\Psi(\mathbf{x},\phi,t)$ is the classical version of Eq. (III) with the substitution
$i\partial_{t}\Psi(\mathbf{x},t)\to i\omega\partial_{\phi}\Psi(\mathbf{x},\phi,t)+i\partial_{t}\Psi(\mathbf{x},\phi,t).$ (128)
The dynamics of the field, accounted by the term $i\partial_{t}\Psi(\mathbf{x},t)$, is then split into the periodic contribution (either determined by the periodicity of the spontaneous Floquet state or by the external driving), parametrized by the phase $\phi$ and encoded in the term $i\omega\partial_{\phi}\Psi(\mathbf{x},\phi,t)$, and the nonperiodic contribution, parametrized by the time $t$ and encoded in the term $i\partial_{t}\Psi(\mathbf{x},\phi,t)$. Once the dynamics is solved for $\Psi(\mathbf{x},\phi,t)$, the physical field is obtained as $\Psi(\mathbf{x},t)=\Psi(\mathbf{x},\phi_{0}+\omega t,t)$, with $\phi_{0}$ some initial phase. When the system is driven, this phase is locked by the actual time-dependence of the Hamiltonian, $H(t)=H(\phi_{0}+\omega t)$, which we can conventionally set to $\phi_{0}=0$. However, for spontaneous Floquet states, $\phi_{0}$ is a free parameter, reflecting the time-independence of the original Hamiltonian.
Intuitively, we can visualize the $(t,\phi)$ space as a cylinder where $t$ parametrizes its longitudinal axis and $\phi$ its polar angle, Fig. 1. The field $\Psi(\mathbf{x},\phi,t)$ is defined over the whole surface of the cylinder, while the actual physical field $\Psi(\mathbf{x},t)$ is evaluated solely along the spiral trajectory $(t,\phi_{0}+\omega t)$ (solid black line).
Regarding the conserved charges, the action is the time integral of the Lagrangian (127), to which we apply Noether Theorem. The extra variable $\phi$ simply adds an angular average to Eq. (12). In particular, the usual charges $\mathbf{Q}$ arising from the continuous symmetries of the Hamiltonian are readily retrieved as
@@
the scalar product in the extended Hilbert space.
@@
with $T_{\phi}=i\partial_{\phi}$ the generator of phase translations. The operator $T_{\phi}$ is Hermitian under the extended scalar product (VI.1), and commutes with the rest of generators, $[T_{\phi},\mathbf{T}]=0$. Finally, invariance under time translations $t\to t+t_{0}$ results in the conservation of the Floquet energy
$I\equiv E-\omega F.$ (132)
Nevertheless, while $I$ is conserved for both driven and spontaneous Floquet states, the conservation of the Floquet charge $F$ is a genuine feature of a spontaneous Floquet state since the Hamiltonian is $\phi$-dependent for driven Floquet systems.
We note that the $(t,\phi)$ formalism can be applied not only to both spontaneous and conventional Floquet systems, but to any many-body system governed by a second-quantization Hamiltonian of the general form (III), including both bosons and fermions, since the dynamics of the field operator then admits a Lagrangian description, as discussed at the beginning of Sec. III.
@@ VI.2 Spontaneous Floquet states in the generalized Gibbs ensemble: Floquet thermodynamics
We invoke again the generalized Gibbs ensemble to describe spontaneous symmetry breaking within the $(t,\phi)$ formalism, replacing the Hamiltonian $H$ by the generalized Gibbs Hamiltonian $K$ in Eq. (127), in analogy with the constrained Lagrangian (20). As a result, the Floquet energy $I$ is replaced by the generalized Floquet-Gibbs energy
$\Lambda\equiv K-\omega F=I-\boldsymbol{\lambda}\cdot\mathbf{Q}=H-\lambda^{\alpha}Q_{\alpha},$ (133)
which is the analogue of the generalized Gibbs energy $K$ for Floquet systems. Its expression allows us to identify $\lambda^{\phi}=\omega$ as the Lagrange multiplier of the Floquet charge.
In the case of a condensate, the generalized Floquet-Gibbs energy $\Lambda$ is the thermodynamic potential that is extremized by stationary solutions $\Psi_{0}(\mathbf{x},\phi)$, satisfying the Floquet-GPG (FGPG) equation
@@
When we insert the FGPG wavefunction in the actual time-dependent GP equation (22) by taking $\Psi(\mathbf{x},t=0)=\Psi_{0}(\mathbf{x},\phi_{0})$, we find
$\Psi(\mathbf{x},t)=e^{-i\lambda^{\alpha}T_{\alpha}t}\Psi_{0}(\mathbf{x},\phi_{0})=e^{-i\boldsymbol{\lambda}\cdot\mathbf{T}t}\Psi_{0}(\mathbf{x},\phi_{0}+\omega t).$ (135)
Thus, Floquet states are described by stationary FGPG solutions with the phase fixed to $\phi=\phi_{0}$, which behaves as a global time origin $t_{0}=\phi_{0}/\omega$. In a spontaneous Floquet state, this time origin can be chosen freely due to the time-independence of the Hamiltonian. In a conventional Floquet state, $t_{0}$ is fixed by the external driving, as discussed after Eq. (128). On the other hand, as expected from thermodynamic considerations, $\omega$ plays the predicted role for the Lagrange multiplier associated to the Floquet charge, acting as the velocity of phase translations. In fact, the definition modulo $\omega$ of the chemical potential $\mu$ for Floquet states can be also explained from Thermodynamics as follows. The transformation $\Psi_{0}(x,\phi)\to\Psi_{0}(x,\phi)e^{-in\phi}$, $n\in\mathbb{Z}$, also yields a solution of the FGPG equation. This leads to $F\to F+nN$, so $\Lambda\to\Lambda-n\omega N$, and hence $\lambda^{\theta}=\mu$ changes as $\mu\to\mu+n\omega$.
@@
we retrieve the generalization of the first principle of Thermodynamics
@@
and the Gibbs-Duhem relation (32) for spontaneous Floquet states.
The upshot of the above discussion is that Floquet states allow for a thermodynamical description completely analogous to that of stationary states, which we will denote as Floquet thermodynamics. Indeed, the first principle of Thermodynamics for spontaneous Floquet states (137) suggests a fundamental distinction between spontaneous and driven Floquet systems: spontaneous systems conserve the charge $F$ and the frequency $\omega$ is determined by the equation of state, while driven systems externally impose the frequency $\omega$ so then $F$ is not conserved. Thus, we can regard spontaneous Floquet systems as “isofloquetic”, where the macroscopic Floquet charge is conserved, in analogy with isentropic or isochoric systems, where the entropy or the volume is conserved. Conversely, conventional Floquet systems are “isoperiodic”, where the external driving fixes the oscillation period, in analogy with isothermal or isobaric systems, where the external environment fixes the temperature or the pressure. Based on this analogy, the Floquet energy $I=E-\omega F$ can be identified as the Floquet enthalpy, the natural thermodynamic potential for isoperiodic systems, where it is conserved and satisfies
@@
VI.3 Quantization
The $(t,\phi)$ formalism emerges in the quantum theory when considering once more the time-independent Heisenberg equation (62). We now perform a unitary transformation to the field operator along the lines of Eq. (135),
$\hat{\Psi}(\mathbf{x},t)=e^{-i\boldsymbol{\lambda}\cdot\mathbf{T}t}\hat{\Psi}(\mathbf{x},\phi+\omega t)=e^{-i\lambda^{\alpha}T_{\alpha}t}\hat{\Psi}(\mathbf{x},\phi,t).$ (139)
The resulting equation of motion for the $(t,\phi)$ field operator reads
$i\hbar\partial_{t}\hat{\Psi}(\mathbf{x},\phi,t)=\left[-\frac{\nabla^{2}}{2}+V(\mathbf{x})+\hat{\Psi}^{\dagger}(\mathbf{x},\phi,t)\hat{\Psi}(\mathbf{x},\phi,t)-\lambda^{\alpha}T_{\alpha}\right]\hat{\Psi}(\mathbf{x},\phi,t),$ (140)
which is the quantum version of the equation of motion that results from the Lagrangian (127). A similar equation can be derived for a conventional Floquet system, where the frequency is fixed by the external driving provided by the periodic Hamiltonian $\hat{H}(\phi)$. Once the dynamics in the $(t,\phi)$ formalism is derived, the actual field operator is retrieved as $\hat{\Psi}(\mathbf{x},t)=e^{-i\boldsymbol{\lambda}\cdot\mathbf{T}t}\hat{\Psi}(\mathbf{x},\phi_{0}+\omega t,t)$. We stress that this equation of motion is exact because it is just a reparametrization of the standard Heisenberg equation of motion. Consequently, the $(t,\phi)$ formalism can be applied along with the Floquet thermodynamics to an arbitrary many-body system.
In the specific case of a condensate, after expanding the field operator around the spontaneous Floquet state, $\hat{\Psi}(\mathbf{x},\phi,t)=\Psi_{0}(\mathbf{x},\phi)+\hat{\varphi}(\mathbf{x},\phi,t)$, we obtain at linear order from Eq. (140) that
@@ VI.3 Quantization
The Goldstone-Gibbs modes $z_{a}(\mathbf{x},\phi)$ are given here in terms of $\partial_{a}\Psi_{0}(\mathbf{x},\phi)$, while the Bogoliubov sector is obtained by the adaptation of the Floquet modes of Eq. (121), $z_{\varepsilon,\nu}(\mathbf{x},\phi)\equiv z_{\varepsilon,\nu}(\mathbf{x},\omega t)$. In order to have a complete orthonormal basis in the extended Hilbert space, we promote each mode $z_{n}(\mathbf{x},\phi)$ to a whole set
$z_{n,m}(\mathbf{x},\phi)\equiv z_{n}(\mathbf{x},\phi)e^{im\phi},~{}m\in\mathbb{Z}.$ (143)
@@
$[M_{0}(\phi)-i\omega\partial_{\phi}]z_{\alpha,m}=m\omega z_{\alpha,m},$
$[M_{0}(\phi)-i\omega\partial_{\phi}]z_{A,m}=m\omega z_{A,m}+i\partial_{A}\lambda^{\beta}z_{\beta,m},$ (144)
$[M_{0}(\phi)-i\omega\partial_{\phi}]z_{\varepsilon,\nu,m}=[\varepsilon+m\omega]z_{\varepsilon,\nu,m},$
@@
$(z_{\varepsilon,\nu,m}||z_{\varepsilon^{\prime},\nu^{\prime},m^{\prime}})$ $=$ $\delta_{\varepsilon\varepsilon^{\prime}}\delta_{\nu\nu^{\prime}}\delta_{mm^{\prime}},$
$(z_{a,m}||z_{\varepsilon,\nu,m^{\prime}})$ $=$ $0,$ (145)
@@
Their quantum amplitude is obtained by $\hat{\gamma}_{n,m}(t)=(z_{n,m}||\hat{\Phi}(t))$. From the Fourier relations
@@
$\hat{\Phi}(\mathbf{x},\phi,t)$ $=$ $\sum_{m=-\infty}^{\infty}\hat{X}_{m}^{\alpha}(t)z_{\alpha,m}(\mathbf{x},\phi)+\hat{P}_{m}^{A}(t)z_{A,m}(\mathbf{x},\phi)+\sum_{m=-\infty}^{\infty}\sum_{\varepsilon,\nu}\hat{\gamma}_{\varepsilon,\nu,m}(t)z_{\varepsilon,\nu,m}(\mathbf{x},\phi)+\hat{\gamma}^{\dagger}_{\varepsilon,\nu.m}(t)\bar{z}_{\varepsilon,\nu,m}(\mathbf{x},\phi)$ (147)
$=$ $\hat{X}^{\alpha}(\phi,t)z_{\alpha}(\mathbf{x},\phi)+\hat{P}^{A}(\phi,t)z_{A}(\mathbf{x},\phi)+\sum_{\varepsilon,\nu}\hat{\gamma}_{\varepsilon,\nu}(\phi,t)z_{\varepsilon,\nu}(\mathbf{x},\phi)+\hat{\gamma}^{\dagger}_{\varepsilon,\nu}(\phi,t)\bar{z}_{\varepsilon,\nu}(\mathbf{x},\phi).$
The dynamics for $\hat{\Phi}(\mathbf{x},\phi,t)$ is readily obtained by the relation $i\partial_{t}\hat{\gamma}_{n,m}(t)=([M_{0}(\phi)-i\omega\partial_{\phi}]z_{n,m}||\hat{\Phi}(t))$, finding:
$\hat{\gamma}_{\varepsilon,\nu,m}(t)=\hat{\gamma}_{\varepsilon,\nu,m}e^{-i[\varepsilon+m\omega]t},~{}~{}~{}\hat{P}_{m}^{A}(t)=\hat{P}_{m}^{A}e^{-im\omega t},~{}~{}~{}\hat{X}_{m}^{\alpha}(t)=[\hat{X}_{m}^{\alpha}+\partial_{A}\lambda_{B}\hat{P}_{m}^{B}t]e^{-im\omega t}.$ (148)
When setting $\phi=\phi_{0}+\omega t$, after noticing that the standard time-dependent modes and amplitudes are $z_{n}(\mathbf{x},t)=z_{n}(\mathbf{x},\phi_{0}+\omega t)$ and $\hat{\gamma}_{n}(t)=\hat{\gamma}_{n}(\phi_{0}+\omega t,t)$, we recover the expected time evolution for the field fluctuations:
$\hat{\Phi}(\mathbf{x},t)=\hat{\Phi}(\mathbf{x},\phi_{0}+\omega t,t)=[\hat{X}^{\alpha}+\partial_{A}\lambda_{B}\hat{P}^{B}t]z_{\alpha}(\mathbf{x},t)+\hat{P}^{A}z_{A}(\mathbf{x},t)+\sum_{\varepsilon,\nu}\hat{\gamma}_{\varepsilon,\nu}z_{\varepsilon,\nu}(\mathbf{x},t)e^{-i\varepsilon t}+\hat{\gamma}^{\dagger}_{\varepsilon,\nu}\bar{z}_{\varepsilon,\nu}(\mathbf{x},t)e^{i\varepsilon t}.$ (149)
The crucial difference with respect to Eq. (VI) is that the Gibbs modes there encode a non-trivial trivial dependence in $z_{\phi}$ since
@@
When this dependence is taken into account, as well as the thermodynamic relation between $K$ and $F$, after some tedious algebra one arrives at the correct expansion (149) from the original Floquet modes of Eqs. (120), (121). We note that the above expansion is also valid for driven Floquet systems by excluding the Goldstone-Gibbs pair $\{z_{\phi},z_{F}\}$ and using $\lambda_{A}=\partial_{A}I$ instead of $\lambda_{A}=\partial_{A}E$.